in two sphers radius r1 and r2 respectivly where r1>r2 we gave them q C charge to each what is the potential diffrence of them

Grade:12

2 Answers

rishabh
10 years ago

kq/r1-kq/r2

 

 

approve by clicking yes

Sumit Majumdar IIT Delhi
10 years ago

Hello student,

The individual potentials of the spheres are as follows:

V1=(1/4∏ε0)(q/r1) and V2=(1/4∏ε0)(q/r2).

Since, the spheres arent connected, so there is no distribution of potentials and hence the potential difference between the spheres would be equal to:

V2-V1=(1/4∏ε0)(q/r2)-(1/4∏ε0)(q/r1)=(1/4∏ε0)q(1/r2-1/r1)

Thanks & Regards
Sumit Majumdar, 
askIITians Faculty
Ph.D, IIT Delhi